OpenStudy (ali2x2):

2+2 = ? D:

OpenStudy (anonymous):

hi, it's 4 :)

Mehek (mehek14):

for real?

OpenStudy (republican31):

2+2=I should click "Report Abuse".

OpenStudy (ali2x2):

answar choices are 21 3 4 22

OpenStudy (ali2x2):

repub its just a joke and its not aganist the rules

Mehek (mehek14):

22

Mehek (mehek14):

2 + 2 = 22

Mehek (mehek14):

:D

OpenStudy (ali2x2):

OMGOMGOM THANK U U SO SMART

Mehek (mehek14):

XD

OpenStudy (ali2x2):

XD

OpenStudy (republican31):

Actually, according to The Fairly Odd Parents on Nickelodeon when I was little, it can be 3 different answers depending on which episode.

Mehek (mehek14):

I can understand 4 and 22

OpenStudy (republican31):

When a certain wish was made, it was known: 2 + 2 = fish. According to Stephen Hawking on the show, it also normally equaled 5. However a 5th grade math teacher was later able to prove him wrong, and that it was actually 6.

OpenStudy (republican31):

Have fun, sir.

Mehek (mehek14):